OpenStudy (anonymous):

at time t=T/4 H/H max is ( where T is the total time of flight h max is the maximum height and H is the height at time t of a projectile launched from the ground)

OpenStudy (anonymous):

did any one get this???

OpenStudy (anonymous):

I'm still solving it

OpenStudy (anonymous):

ok try ur best............

OpenStudy (anonymous):

I got 7/16, if that is right answer I'll give you the whole solution.

OpenStudy (anonymous):

the answer is 3/4

OpenStudy (anonymous):

let's assume that v0y=v0 t1=T/2 time to reach the Hmax height \[H_{max}=v_{0}t_{1}-gt_{1}^2/2\] [1] \[v=v_{0}-gt\] \[v(t_{1})=0=v_{0}-gt_{1}\] \[v_{0}=gt_{1}\] \[g=v_{0}/t_{1}\] using [1], \[H_{max}=v_{0}t_{1}-(v_{0}/t_{1})t_{1}^2/2=v_{0}t_{1}/2=v_{0}T/4\] \[H_{max}=v_{0}T/4\] \[H(t) = v_{0}t-(v_{0}/t_{1})t^2/2\] \[H(T/4)=v_{0}T/4-(v_{0}/t_{1})T/4^2/2=v_{0}T/4-(2v_{0}/T)T/4^2/2=\] \[v_{0}T/4-v_{0}/16= (3/16)v_{0} T\] \[H=(3/16)v_{0} T\] now take the ratio

OpenStudy (anonymous):

THANXXXXX LAST ONE MORE QUESTION PLZZ SEE IT